# Improving the Euclidean Diffusion Generation of Manifold Data by Mitigating Score Function Singularity

### v1

@@ Abstract
Euclidean diffusion models have achieved remarkable success in generative modeling across diverse domains, and they have been extended to manifold case in recent advances. Instead of explicitly utilizing the structure of special manifolds as studied in previous works, we investigate direct sampling of the Euclidean diffusion models for general manifold-constrained data in this paper. We reveal the multiscale singularity of the score function in the embedded space of manifold, which hinders the accuracy of diffusion-generated samples. We then present an elaborate theoretical analysis of the singularity structure of the score function by separating it along the tangential and normal directions of the manifold. To mitigate the singularity and improve the sampling accuracy, we propose two novel methods: (1) Niso-DM, which introduces non-isotropic noise along the normal direction to reduce scale discrepancies, and (2) Tango-DM, which trains only the tangential component of the score function using a tangential-only loss function. Numerical experiments demonstrate that our methods achieve superior performance on distributions over various manifolds with complex geometries.
@@
1 Introduction
Recent advances in diffusion models [28, 9, 29, 31] have demonstrated remarkable success in generative modeling in various domains, including image generation [10], audio synthesis [3], molecule generation [11], and other applications [36, 16]. These models operate through a forward process, which gradually perturbs data into noise, and a reverse process, which reconstructs the data from noise.
Beyond Euclidean space, many scientific fields involve data distributions constrained to Riemannian manifolds. Examples include geographical sciences [22] using spheres, protein structures [34] and robotic movements [27] with $\mathrm{SE}(3)$ and $\mathrm{SO}(3)$, lattice quantum chromodynamics [20] with $\mathrm{SU}(3)$, 3D computer graphics [12] with triangular meshes, and cell development [15] with the Poincaré disk. To model such data, recent works [7, 20, 13, 37] have extended diffusion models to Riemannian manifolds, achieving notable progress by leveraging manifold geometry and curvature information. Nonetheless, these methods face challenges due to the cost of simulating diffusion and obtaining accurate geometric information like the heat kernel.
On the other hand, a natural idea is to directly apply diffusion models designed for Euclidean space to manifold-structured data. However, this approach encounters a fundamental challenge: the singularity of the score function. Prior studies under the manifold hypothesis [23, 21, 6] have identified that in the reverse process, the norm of the score function grows explosively. Recent works [21, 23, 2] have theoretically analyzed this singularity under the Ornstein-Uhlenbeck process, providing asymptotic bounds on the score function and introducing structural assumptions on the data distribution. In this paper, we further enhance these insights and propose methods for mitigating the singularity of the score function in the Euclidean space.
Our work. We consider a probability distribution on a known $d$-dimensional submanifold $\mathcal{M}$ embedded in $\mathbb{R}^{n}$. Given samples from this distribution, our objective is to generate new samples by directly applying Euclidean diffusion models. First, under the Variance Exploding Stochastic Differential Equations (VESDE) framework with such a manifold setting, we theoretically demonstrate that the perturbed score function exhibits distinct scales along the tangential and normal components. To address this issue, we propose the following two methods to improve diffusion models (DM):
@@ 1 Introduction
    Niso-DM: Perturbs data with non-isotropic noise by introducing additional noise along the normal direction during the forward diffusion process. This reduces the scale discrepancy between the tangential and normal directions, making the relaxed score function easier to approximate.
@@
    Tango-DM: Trains only the tangential component of the score function using a tangential-only loss function. By skipping the training of the normal component, this method avoids the learning difficulties caused by the multiscale issue.
@@
  • We propose two novel methods, Niso-DM and Tango-DM, making the relaxed score function easier to approximate. We also give theoretical analysis of the Niso-DM. Empirically, our methods demonstrate superior performance on distributions across various complex and non-trivial manifolds. This is in sharp contrast with previous methods which can only handle special manifolds with particular analytical structure.
@@ 2.1 Diffusion models
Diffusion models can be formulated using stochastic differential equations (SDEs). Specifically, data perturbation is modeled by the following forward SDE:
@@
where $f$ and $g$ are fixed drift and diffusion coefficients, respectively. The time-dependent distribution $p_{t}$, driven by the stochastic process $X_{t}$, converges to a limiting distribution $\mathcal{N}(0,\sigma_{T}^{2}I)$ as $t\to+\infty$. By carefully selecting $f$ and $g$, this distribution can be approximated as $p_{T}\approx\mathcal{N}(0,\sigma_{T}^{2}I)$. The corresponding reverse-time SDE, which transports $p_{T}$ back to the initial probability density function (pdf) $p_{0}$, is given by:
@@
where $\overline{W}_{t}$ is a standard Brownian motion in reverse time.
@@
$\mathcal{L}_{\text{quad}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1)}\lambda_{t}\ell_{\text{quad}}(t,\theta),$ (3)
$\ell_{\text{quad}}(t,\theta)=\mathbb{E}_{x_{t}\sim p_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{t}(x_{t})\|^{2},$ (4)
is intractable. Instead, score matching techniques such as denoising score matching [33] are commonly used, formulated as:
$\ell_{\text{dsm}}(t,\theta)=\mathbb{E}_{x,x_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{t}(x_{t}|x)\|^{2}.$ (5)
To utilize the above loss, the computability of $p_{t}(x_{t}|x)$ requires specific SDEs, such as the following VPSDE and VESDE introduced by [31].
@@
    $\mathrm{d}X_{t}=\frac{1}{2}\beta(t)X_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}W_{t}.$ (6)
@@
    where $\sigma_{t}$ is a predefined noise scale, which is commonly chosen as $\sigma_{t}=\sigma_{\min}(\sigma_{\max}/\sigma_{\min})^{t/T}$. In this case, $p_{t}(x_{t}|x)$ follows a Gaussian distribution: $\mathcal{N}(x_{t}|x,\sigma_{t}^{2}I)$.
Choice of SDE in this work. In this study, we choose VESDE for its advantageous properties. Under VESDE, the mean of the perturbed distribution remains unchanged, preserving the original manifold structure. At any time, the perturbed distribution $p_{t}(x)$ can be viewed as points on the original manifold $\mathcal{M}$ with added noise. This simplifies theoretical analysis and avoids the complications introduced by the evolving manifold in VPSDE. In fact, VPSDE involves an evolving manifold over time, defined as $\mathcal{M}_{t}=\{e^{-\frac{1}{2}\int_{0}^{t}\beta(s)\mathrm{d}s}x\mid x\in\mathcal{M}\}$ as the mean with noise perturbation.
@@ 2.2 Riemannian manifolds and notation
In this paper, the manifold $\mathcal{M}$ is viewed as =$\mathcal{M}=\{x\in\mathbb{R}^{n}|\xi(x)=0\}$ embedded in $\mathbb{R}^{n}$, $\xi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n-d}$ is a known function, and we assume that $\nabla\xi(x)\in\mathbb{R}^{n\times(n-d)}$ has always full column rank. The target distribution is formulated as $p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$, where $p_{0}(x)$ is an unknown density and $\mathrm{d}\sigma_{\mathcal{M}}(x)$ represents the volume form of $\mathcal{M}$. The unit normal vector at a point $x\in\mathcal{M}$ is given by $N(x)=\nabla\xi(x)\left(\nabla\xi(x)^{T}\nabla\xi(x)\right)^{-\frac{1}{2}}$. For $x\in\mathcal{M}$, the Riemannian score function is defined as $\nabla_{x}^{\mathcal{M}}\log p_{0}(x)=P(x)\nabla_{x}\log p_{0}(x)$, where $P(x)$ is the projection matrix onto the tangent space of $\mathcal{M}$. The projection matrix $P(x)$ is computed as
@@ 3 Scale discrepancy of score functions
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=\int_{\mathcal{M}}p_{0}(x)(2\pi\sigma^{2})^{-\frac{n}{2}}e^{-\frac{|x-\tilde{x}|^{2}}{2\sigma^{2}}}\mathrm{d}\sigma_{\mathcal{M}}(x),$ (9)
where $\tilde{x}$ is a variable in the embedded space $\mathbb{R}^{n}$, $p_{\sigma}(\tilde{x}|x)$ denotes the isotropic perturbation pdf $\mathcal{N}(x,\sigma^{2}I)$, $p_{\sigma}(\tilde{x})$ is the perturbed pdf defined on $\mathbb{R}^{n}$, and $p_{0}(x)$ is only defined on $\mathcal{M}$. Hereafter, the subscript of $p$ denotes the noise scale $\sigma$ instead of the time $t$.
When the noise scale $\sigma$ is small, the perturbed distribution becomes tightly concentrated around its mean, resulting in a steep gradient landscape for $-\log p_{\sigma}(\tilde{x})$. As $\sigma$ approaches zero, this steepness appears as sharp variations and introduces multiscale challenges. In Theorem 3.1, we provide a rigorous and refined results on this singularity compared to previous works. Eq. (10) shows that the score function explodes entirely due to the normal direction to the manifold, represented by $x^{*}-\tilde{x}$. After subtracting this term, the remaining components are of order $O(1)$. Furthermore, Equations (11) and (12) establish a novel connection between the Riemannian score function and the score perturbed function in the ambient space for points on the manifold, which has not been considered in previous works. The proof is provided in Appendix A.
@@
Let $P(x)\in\mathbb{R}^{n\times n}$ denote the projection matrix at $x\in\mathcal{M}$. Assume that $\mathbb{E}_{p_{0}}\|x\|<+\infty$ and $M_{1}=\sup_{x\in\mathcal{M}}\max_{1\leq i,j,j^{\prime}\leq n}\left|\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}(x)\right|<+\infty$. The following two asymptotic expansions for $p_{\sigma}(\tilde{x})$ defined in (9) hold:
@@
    $\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\nabla_{\tilde{x}}^{\mathcal{M}}\log p_{0}(\tilde{x})-\frac{1}{2}\sum_{j,j^{\prime}=1}^{n}\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}(\tilde{x})P_{jj^{\prime}}(\tilde{x})+O(\sigma),$ (11)
    $\nabla_{\tilde{x}}^{\mathcal{M}}\log p_{\sigma}(\tilde{x})=\nabla_{\tilde{x}}^{\mathcal{M}}\log p_{0}(\tilde{x})+O(\sigma).$ (12)
Next, based on Theorem 3.1, we analyze the scale discrepancy of the score function between its tangential and normal components. To achieve this, we first extend the projection matrix $P$ from $\mathcal{M}$ to the entire space $\mathbb{R}^{n}$ continuously. A natural extension is to define $P$ in the same manner as in (8), assuming that $\xi$ is well-defined and smooth on $\mathbb{R}^{n}$. With this extended projection matrix, we further define the tangential and normal components of the score function for any $\tilde{x}\in\mathbb{R}^{n}$ as $P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})$ and $P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})$, respectively, where we denote $P^{\perp}(\tilde{x})=I-P(\tilde{x})$. Using this extended definition, we then decompose the quadratic loss $\ell_{\text{quad}}$ into two parts: $\ell_{\text{quad}}^{\parallel}$ and $\ell_{\text{quad}}^{\perp}$, corresponding to the tangential and normal components, respectively.
$\ell_{\text{quad}}=$ $\mathbb{E}_{\tilde{x}\sim p_{\sigma}(\tilde{x})}\|s_{\theta}(\tilde{x},t)-\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$ (13)
$=$ $\mathbb{E}_{\tilde{x}\sim p_{\sigma}(\tilde{x})}\|P(\tilde{x})s_{\theta}(\tilde{x},t)-P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$
$+\mathbb{E}_{\tilde{x}\sim p_{\sigma}(\tilde{x})}\|P^{\perp}(\tilde{x})s_{\theta}(\tilde{x},t)-P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$
@@
Using Eq. (10) and the facts that $P(x^{*})(x^{*}-\tilde{x})=0$, $P(\tilde{x})=P(x^{*})+O(\tilde{x}-x^{*})$ and $\mathbb{E}_{\tilde{x}|x}(x^{*}-\tilde{x})=O(\sigma)$, where $\mathbb{E}_{\tilde{x}|x}$ denotes the expectation with respect to $p_{\sigma}(\tilde{x}|x)$, the two terms being approximated in (13) have scales of $O(1)$ and $O(1/\sigma)$, respectively, by the following two estimates:
$\mathbb{E}_{\tilde{x}|x}P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\mathbb{E}_{\tilde{x}|x}\left(P(x^{*})+O(\tilde{x}-x^{*})\right)\Big{(}\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\Big{)}=O(1),$ (14)
$\mathbb{E}_{\tilde{x}|x}P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})=\mathbb{E}_{\tilde{x}|x}\left(P^{\perp}(x^{*})+O(\tilde{x}-x^{*})\right)\Big{(}\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\Big{)}=O\Big{(}\frac{1}{\sigma}\Big{)}.$ (15)
This multiscale singularity of the loss formulation poses significant challenges during training. Since the training loss (5) is essentially a quadratic loss (4), it inherently prioritizes fitting larger-scale features of the score, which are aligned with the normal component in our settings. This is because quadratic loss minimizes the average of squared errors, making it disproportionately sensitive to larger errors that predominantly occur in higher-scale directions. In the context of this work, the normal component pulls samples back onto the manifold, whereas the tangential component, which is more critical, captures the data distribution on the manifold. As a result, the model fails to adequately capture finer details of the data distribution, ultimately reducing the accuracy of the generated distribution. In other words, the trained model primarily captures the manifold itself rather than the distribution on it, leading to a phenomenon known as manifold overfitting [18].
To address these limitations, specific methods are required during the model training phase to ensure training stability and accurately capture the intrinsic data distribution on the manifold. Based on this perspective, we propose the following two solutions: First, the scale discrepancy between the tangential and normal components can be reduced by modifying the structure of the noise. Second, instead of explicitly considering the normal component of the score function, its role in projecting samples back onto the manifold during generation is delegated to a dedicated projection operator.

### v3

@@ Abstract
Euclidean diffusion models have achieved remarkable success in generative modeling across diverse domains, and they have been extended to manifold cases in recent advances. Instead of explicitly utilizing the structure of special manifolds as studied in previous works, in this paper we investigate direct sampling of the Euclidean diffusion models for general manifold-structured data. We reveal the multiscale singularity of the score function in the ambient space, which hinders the accuracy of diffusion-generated samples. We then present an elaborate theoretical analysis of the singularity structure of the score function by decomposing it along the tangential and normal directions of the manifold. To mitigate the singularity and improve the sampling accuracy, we propose two novel methods: (1) Niso-DM, which reduces the scale discrepancies in the score function by utilizing a non-isotropic noise, and (2) Tango-DM, which trains only the tangential component of the score function using a tangential-only loss function. Numerical experiments demonstrate that our methods achieve superior performance on distributions over various manifolds with complex geometries.
1 Introduction
@@
Diffusion models [35, 11, 36, 38] have demonstrated remarkable success in generative modeling in various domains, including image generation [12], audio synthesis [5], molecule generation [13], and other applications [48, 22]. These models operate through a forward process, which gradually perturbs data into noise, and a reverse process, which reconstructs the data from noise.
Beyond Euclidean space, many scientific fields involve data distributions constrained on Riemannian manifolds. Examples include geographical sciences [28] using spheres, protein structures [45] and robotic movements [34] with $\mathrm{SE}(3)$ and $\mathrm{SO}(3)$, lattice quantum chromodynamics [26] with $\mathrm{SU}(3)$, 3D computer graphics [14] with triangular meshes, and cell development [21] with the Poincaré disk. To model such data, recent works [8, 26, 17, 49] have extended diffusion models to Riemannian manifolds, achieving notable progress by leveraging manifold geometry. Nonetheless, these methods face challenges due to the computational complexity of simulating diffusion and obtaining accurate geometric information like the heat kernel.
On the other hand, a natural idea is to directly apply diffusion models designed for Euclidean space to manifold-structured data. However, this approach encounters a fundamental challenge due to the singularity of the score function. Prior studies [29, 27, 3] have identified that under the manifold hypothesis, the norm of the score function explodes as the time in the reverse process approaches zero. Recent works [27, 29, 4] have theoretically analyzed this singularity under the Ornstein-Uhlenbeck process, providing asymptotic bounds on the score function and introducing structural assumptions on the data distribution. In this paper, we further enhance these insights and propose methods for mitigating the singularity of the score function under the manifold hypothesis.
Our work. We consider a probability distribution on a known $d$-dimensional submanifold $\mathcal{M}$ embedded in $\mathbb{R}^{n}$, where $d<n$. Given samples from this distribution, our objective is to generate new samples by directly applying Euclidean diffusion models. Within the framework of Variance Exploding Stochastic Differential Equations (VESDE) with such a manifold setting, we theoretically demonstrate that the perturbed score function exhibits distinct scales along the tangential and normal components. To address this issue, we propose the following two methods to improve diffusion models (DM):
@@
    Niso-DM: Perturb data with non-isotropic noise by introducing additional noise along the normal direction during the forward diffusion process. This reduces the scale discrepancy in the score function between the tangential and normal directions, making the relaxed score function easier to approximate.
@@
    Tango-DM: Train only the tangential component of the score function using a tangential-only loss function. By bypassing the training of the normal component, this method overcomes the learning difficulties caused by the multiscale issue.
@@
  • We propose two novel methods, Niso-DM and Tango-DM, making the relaxed score function easier to approximate. We also give a theoretical analysis of Niso-DM. Empirically, our methods demonstrate superior performance on distributions across various complex and non-trivial manifolds. This is in sharp contrast with previous methods that can only handle manifolds with special analytical structure.
@@ 2.1 Diffusion models
Diffusion models can be formulated using stochastic differential equations (SDEs) [38]. Specifically, the data perturbation process is modeled by the following forward SDE:
@@
where $f$ and $g$ are fixed drift and diffusion coefficients, respectively. By carefully selecting $f$ and $g$, the distribution of $X_{T}$, with the probability density function (pdf) $p_{T}$, can be well approximated by a Gaussian distribution $\mathcal{N}(0,\sigma_{T}^{2}I)$, where $\sigma_{T}^{2}$ represents the variance of the Gaussian distribution. The corresponding reverse-time SDE, which transports $p_{T}$ back to the initial pdf $p_{0}$, is given by:
@@
where $\overline{W}_{t}$ is a standard Brownian motion in reverse time and $p_{t}$ is the time-dependent distribution driven by the stochastic process $X_{t}$.
@@
$\mathcal{L}_{\text{quad}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1)}[\lambda_{t}\ell_{\text{quad}}(t,\theta)],$ (3)
$\ell_{\text{quad}}(t,\theta)=\mathbb{E}_{x_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{t}(x_{t})\|^{2},$ (4)
is intractable. Instead, score matching techniques such as denoising score matching [43] are commonly used, which learn the score with the loss
$\ell_{\text{dsm}}(t,\theta)=\mathbb{E}_{x,x_{t}}\|s_{\theta}(x_{t},t)-\nabla_{x_{t}}\log p_{t}(x_{t}|x)\|^{2},$ (5)
where $p_{t}(x_{t}|x)$ denotes the probability density of $X_{t}$ in (1) conditioned on the initial state $X_{0}=x$. To utilize (5), the following two specific SDEs introduced in [38] are often considered, in which cases the density $p_{t}(x_{t}|x)$ has a closed-form.
@@
    $\mathrm{d}X_{t}=-\frac{1}{2}\beta(t)X_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}W_{t},$ (6)
    where $\beta(t)>0$.
@@
    where $\sigma_{t}$ is a predefined noise scale, commonly chosen as $\sigma_{t}=\sigma_{\min}(\sigma_{\max}/\sigma_{\min})^{t/T}$ for some $\sigma_{\max}>\sigma_{\min}>0$. In this case, $p_{t}(x_{t}|x)$ follows a Gaussian distribution $\mathcal{N}(x_{t}|x,\sigma_{t}^{2}I)$.
In this study, we choose VESDE for its advantageous properties. Under VESDE, the mean of the perturbed distribution remains unchanged. At any time, the perturbed distribution $p_{t}(x)$ can be viewed as points on the original manifold $\mathcal{M}$ with added noise. This simplifies theoretical analysis and avoids the complications introduced by the evolving manifold in VPSDE. In fact, VPSDE involves an evolving manifold over time, defined as $\mathcal{M}_{t}=\{e^{-\frac{1}{2}\int_{0}^{t}\beta(s)\mathrm{d}s}x\mid x\in\mathcal{M}\}$ as the mean with noise perturbation.
@@ 2.2 Riemannian manifolds and notation
In this paper, the manifold $\mathcal{M}$ is viewed as $\mathcal{M}=\{x\in\mathbb{R}^{n}|\xi(x)=0\}$ embedded in $\mathbb{R}^{n}$, $\xi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n-d}$ is a known function, and we assume that $\nabla\xi(x)\in\mathbb{R}^{n\times(n-d)}$ has full column rank for all $x\in\mathcal{M}$. The target distribution is formulated as $p_{0}(x)\mathrm{d}\sigma_{\mathcal{M}}(x)$, where $p_{0}(x)$ is an unknown density and $\mathrm{d}\sigma_{\mathcal{M}}(x)$ represents the volume form [32] of $\mathcal{M}$. The unit normal vectors at a point $x\in\mathcal{M}$ are given by $N(x)=\nabla\xi(x)\left(\nabla\xi(x)^{T}\nabla\xi(x)\right)^{-\frac{1}{2}}$. For $x\in\mathcal{M}$, the operator $\nabla_{x}^{\mathcal{M}}$ is defined as $\nabla_{x}^{\mathcal{M}}h(x)=P(x)\nabla_{x}h(x)$ for any smooth function $h$, where $P(x)$ is the projection matrix [32] onto the tangent space of $\mathcal{M}$ and is given by
@@ 3 Scale discrepancy of score functions
$p_{\sigma}(\tilde{x})=\int_{\mathcal{M}}p_{0}(x)p_{\sigma}(\tilde{x}|x)\mathrm{d}\sigma_{\mathcal{M}}(x)=(2\pi\sigma^{2})^{-\frac{n}{2}}\int_{\mathcal{M}}p_{0}(x)e^{-\frac{|x-\tilde{x}|^{2}}{2\sigma^{2}}}\mathrm{d}\sigma_{\mathcal{M}}(x),$ (9)
where $\tilde{x}$ is a variable in the embedded space $\mathbb{R}^{n}$, $p_{\sigma}(\tilde{x}|x)$ denotes the pdf $\mathcal{N}(x,\sigma^{2}I)$ of the isotropic perturbtion, and $p_{0}(x)$ is only defined on $\mathcal{M}$. Hereafter, the subscript of $p$ denotes the noise scale $\sigma$ instead of the time $t$.
[…]
@@
[…]
@@
[…]
    where $\frac{\partial P_{\cdot j}}{\partial x_{j^{\prime}}}$ denotes the vector whose $i$th component is $\frac{\partial P_{ij}}{\partial x_{j^{\prime}}}$.
[…]
$\ell_{\text{quad}}=$ $\mathbb{E}_{\tilde{x}}\|s_{\theta}(\tilde{x},t)-\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$ (13)
$=$ $\mathbb{E}_{\tilde{x}}\|P(\tilde{x})s_{\theta}(\tilde{x},t)-P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}+\mathbb{E}_{\tilde{x}}\|P^{\perp}(\tilde{x})s_{\theta}(\tilde{x},t)-P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})\|^{2}$
@@
Using Eq. (10) and the facts that $P(x^{*})(x^{*}-\tilde{x})=0$, $P(\tilde{x})=P(x^{*})+O(\tilde{x}-x^{*})$ and $\mathbb{E}_{\tilde{x}|x}(x^{*}-\tilde{x})=O(\sigma)$, the two terms being approximated in (13) have scales of $O(1)$ and $O(1/\sigma)$, respectively, by the following two estimates:
$\mathbb{E}_{\tilde{x}|x}[P(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})]=\mathbb{E}_{\tilde{x}|x}\left[\left(P(x^{*})+O(\tilde{x}-x^{*})\right)\Big(\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\Big)\right]=O(1),$ (14)
$\mathbb{E}_{\tilde{x}|x}[P^{\perp}(\tilde{x})\nabla_{\tilde{x}}\log p_{\sigma}(\tilde{x})]=\mathbb{E}_{\tilde{x}|x}\left[\left(P^{\perp}(x^{*})+O(\tilde{x}-x^{*})\right)\Big(\frac{x^{*}-\tilde{x}}{\sigma^{2}}+O(1)\Big)\right]=O\Big(\frac{1}{\sigma}\Big),$ (15)
where $\mathbb{E}_{\tilde{x}|x}$ denotes the expectation with respect to $p_{\sigma}(\tilde{x}|x)$.
This multiscale singularity of the loss formulation poses significant challenges during training. In fact, the above analysis also applies to the loss (5), since it differs from (4) only by a constant that is independent of $\theta$. As a result, training with (5) inherently prioritizes fitting larger-scale features of the score, which are aligned with the normal component in our settings. This is because quadratic loss minimizes the average of squared errors, making it disproportionately sensitive to errors that occur in larger-scale directions. In the context of this work, the normal component pulls samples back onto the manifold, whereas the tangential component, which is more critical, captures the data distribution on the manifold. As a result, the model fails to adequately capture finer details of the data distribution, ultimately reducing the accuracy of the generated distribution. In other words, the trained model primarily captures the manifold itself rather than the distribution on it, leading to a phenomenon known as manifold overfitting [24].
To address these limitations, specific methods are required during the model training phase to ensure training stability and accurately capture the intrinsic data distribution on the manifold. Based on this perspective, we propose the following two methods. The first method reduces the scale discrepancy between the tangential and normal components by modifying the structure of the noise. The second bypasses the learning of the normal component of the score function and employs a dedicated projection operator to project samples back onto the manifold.